### v8

# A Unified Graphical Criterion for Characterizing a Linear Causal Interpretation of Partial Regression Coefficients

Abstract
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. However, a general condition has not yet been fully established for settings in which mediators or nonlinear terms are included among the explanatory variables. To address this issue, this paper unifies d-separation, single-door, and back-door criteria and proposes a generalized graphical criterion that enables the identification of causal effects not mediated by other explanatory variables. In other words, for almost all parameter points in the parameter space of linear causal models, this criterion is a necessary and sufficient condition under which partial regression coefficients have a linear causal interpretation. In addition, this paper provides a quantitative characterization of post-treatment bias while controlling for confounding bias. These results apply to linear models represented by acyclic directed mixed graphs, where the models can include nonlinear terms. The decomposition of linear and nonlinear causal effects is also discussed.
1 Introduction
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. Among existing approaches, the back-door criterion (Pearl, 1998) and the conditional ignorability assumption (Rubin, 1974) are often regarded as among the most refined conditions. When either of these conditions and some additional assumptions are satisfied, partial regression coefficients in linear structural equation models (SEMs), as developed by Wright (1921), Haavelmo (1943), Jöreskog (1970), Goldberger (1972), and Bollen (1989), can be interpreted as total effects.
However, the following three issues remain unresolved: (1) the control of specific mediated effects; (2) understanding the mechanism underlying post-treatment bias; and (3) causal interpretation in the presence of nonlinear terms. Issues (1) and (2) are fundamental topics in mediation analysis. Since nonlinear terms can be regarded as mediators, a deeper understanding of how conditioning on mediators affects partial regression coefficients is necessary for capturing complex causal relationships.
In relation to issue (1), the single-door criterion (Pearl, 1998) and the sequential ignorability assumption (Imai et al., 2010b) provide partial answers; nevertheless, these conditions cannot accommodate certain mediated effects along specific causal paths. For example, consider the following linear regression equation at the population level:
$Y$ $=\beta_{Y0\mid M_{1}X}+\beta_{YM_{1}\mid X}M_{1}+\beta_{YX\mid M_{1}}X+\epsilon_{Y\mid M_{1}X},$ (1.1)
where $X$, $M_{1}$, and $Y$ are generated by the linear SEM corresponding to Figure 1.
Figure 1: The causal (path) diagram underlying (1.1)
If we are interested in the mediated effect along the path $X\rightarrow M_{2}\rightarrow Y$ in Figure 1, neither condition clearly indicates whether this effect can be identified via the partial regression coefficient $\beta_{YX\mid M_{1}}$. As it turns out, this coefficient still coincides with the causal effect of interest, implying that the existing conditions can be generalized.
As shown in (1.1), mediation analysis often makes use of post-treatment variables to block certain causal paths. When conditioning on a mediator, one might expect to control only for the mediated effect; however, it may also induce unintended bias. For example, a variable such as $M_{2}$ in Figure 1, which is influenced by multiple variables, is referred to as a collider, and it is well known that conditioning on a collider can introduce such bias. Other structures that give rise to post-treatment bias are also discussed by Elwert and Winship (2014) and Pearl (2015). Nevertheless, the mechanism underlying post-treatment bias has not yet been fully elucidated, even when restricted to linear models.
In regression analysis, nonlinear terms are often included to capture complex relationships. For example, if the data-generating process of $M_{2}$ is given by $M_{2}=XM_{1}$, the directed edges $X\rightarrow M_{2}$ and $M_{1}\rightarrow M_{2}$ in Figure 1 represent nonlinear causal relationships. In such cases, Imai et al. (2010a) and VanderWeele (2013) examine the identification of total causal effects of $X$ on $Y$ using partial regression coefficients. However, these results rely on specific structural assumptions; therefore, it is necessary to discuss this problem in a more generalized setting.
Against this background, we propose a generalized graphical criterion and provide, in a generic sense, a necessary and sufficient condition under which a partial regression coefficient coincides with a causal effect not mediated by other explanatory variables. In addition, we provide a quantitative characterization of post-treatment bias while controlling for confounding bias. This characterization shows that the joint structure of directed paths and back-door paths is the source of this bias, and that there exist structures that cannot be captured by the conventional concept of path blocking. Finally, we show that the above results can be applied to nonlinear data-generating processes and discuss a decomposition of linear and nonlinear causal effects.
The remainder of this paper is organized as follows. Section 2 introduces key terms as preliminaries. Section 3 formulates the problem addressed in this paper and presents the main theorems. This section also discusses the mechanisms underlying post-treatment bias and applications to nonlinear causal models. All proofs are presented in the Appendix.
2 Preliminaries
2.1 Model Setup
Let $\bm{V}=\{X_{1},\dots,X_{N}\}$ be a finite set of random variables with nonzero variances. Note that $\bm{V}$ may include unobservable variables. This paper considers the following model for the data-generating process:
$X_{i}=\sum_{X_{j}\in\bm{PA}(X_{i})}a_{ij}X_{j}+u_{i},\quad X_{i}\in\bm{V},$ (2.1)
where $\bm{PA}(X_{i})\subset\bm{V}\setminus\{X_{i}\}$ denotes the set of immediate causes of $X_{i}$. Here, $a_{ij}$ is a nonzero constant referred to as the path coefficient from $X_{j}$ to $X_{i}$, and $u_{1},\dots,u_{N}$ are error terms representing aggregations of omitted causes. We do not assume any specific distribution for the error terms.
model (2.1) can be rewritten as follows:
$X_{i}=c_{i}+\sum_{X_{j}\in\bm{PA}(X_{i})}a_{ij}X_{j}+\tilde{u}_{i},\quad X_{i}\in\bm{V},$
where $c_{i}=\,\mathrm{E}[u_{i}]$ and $\tilde{u}_{i}=u_{i}-\,\mathrm{E}[u_{i}]$. Moreover, the following vector form is useful:
$\bm{X}=\bm{A}\bm{X}+\bm{u}=\bm{c}+\bm{A}\bm{X}+\tilde{\bm{u}}$, (2.2)
where $\bm{X}=(X_{1},\dots,X_{N})^{\top}$, $\bm{u}=(u_{1},\dots,u_{N})^{\top}$, and $\tilde{\bm{u}}=(\tilde{u}_{1},\dots,\tilde{u}_{N})^{\top}$. Here, $\bm{A}$ is an $N\times N$ matrix, and $\bm{c}$ is an $N\times 1$ vector. We make the following assumptions on (2.2).
Assumption 2.1.
  • (i)
    $\bm{\Sigma}\coloneq\mathop{\rm Var}\nolimits(\bm{u})$ is a positive definite matrix.
  • (ii)
    $\bm{A}$ is a lower triangular matrix with all diagonal elements equal to zero.
Assumption 2.1 ensures that the graph associated with model (2.1) is an acyclic directed mixed graph (ADMG), discussed by (Richardson, 2003). With respect to the coefficient matrix $\bm{A}$, strictly speaking, it is sufficient that $\bm{A}$ can be transformed into the structure described in (ii) by reordering the indices of the random variables in $\bm{V}$. However, note that unless the graph structure is known, graphical criteria are in most cases not applicable. In some cases, the graph structure can be identified using methods such as LiNGAM (Shimizu et al., 2006). When the graph structure is unknown, it is necessary to conduct the analysis under each plausible structure and compare the results.
2.2 Graph Terminology and Notations
An ADMG is a triple $(\bm{V},\bm{E},\bm{B})$, where $\bm{V}$ is a finite set of vertices, $\bm{E}$ is a set of directed edges that construct no directed cycles, and $\bm{B}$ is a set of bidirected edges. For vertices $a,b\in\bm{V}$, if they are connected by either a directed edge (denoted by $a\rightarrow b$) or a bidirected edge (denoted by $a\leftrightarrow b$), then $a$ and $b$ are said to be adjacent. If there is a directed edge from $a$ to $b$, then $a$ is called a parent of $b$, and $b$ is called a child of $a$. A path is a sequence of distinct vertices such that consecutive vertices are adjacent. A directed path from $a$ to $b$ (denoted by $a\rightarrow\rightarrow b$) is a path along which all edges are directed and oriented from $a$ toward $b$. A directed cycle is a concatenation of a directed path $a\rightarrow\rightarrow b$ and a directed edge $b\rightarrow a$. If there is a directed path from $a$ to $b$, then $a$ is called an ancestor of $b$, and $b$ is called a descendant of $a$. We denote the sets of ancestors and descendants of $a$ by $\bm{AN}(a)$ and $\bm{DE}(a)$, respectively. A back-door path from $a$ to $b$ (denoted by $a\dashleftarrow b$) is a path from $a$ to $b$ that begins with either a bidirected edge or a directed edge pointing into $a$. A v-structure is a path of the form $a\rightarrow c\leftarrow b$, $a\rightarrow c\leftrightarrow b$, $a\leftrightarrow c\leftarrow b$, or $a\leftrightarrow c\leftrightarrow b$, where $c$ is called a collider.
2.3 Causal Path Diagram and the Selective-Door Criterion
The discussion in this paper is based on linear models, and the interpretation of bidirected edges is therefore slightly different from that in standard causal diagrams (Pearl, 2009). Accordingly, we distinguish our framework by using the term causal path diagram.
Definition 2.2 (Causal path diagram).
Consider model (2.1). Let $\bm{V}$ denote the set of vertices, and let $\bm{E}$ denote the set of directed edges such that, for each $X_{i}\in\bm{V}$, the set $\bm{PA}(X_{i})$ coincides with the parent set of $X_{i}$. In addition, let $\bm{B}$ denote the set of bidirected edges given by
[…]
Then, the triple $(\bm{V},\bm{E},\bm{B})$ is called the causal path diagram associated with model (2.1).
Next, we introduce the following concepts to visually identify the conditional relationships between variables.
Definition 2.3 (Blocking).
Let $\bm{Z}$ be a set such that $\bm{Z}\cap\{X_{i},X_{j}\}=\emptyset$. The set $\bm{Z}$ is said to block a path $P$ between $X_{i}$ and $X_{j}$ if it satisfies one of the following two conditions; otherwise, $\bm{Z}$ is said to open the path:
  • (i)
    $\bm{Z}$ contains a vertex on $P$ that is not a collider in any v-structure along $P$; or
  • (ii)
    For a v-structure along $P$, neither the collider nor any of its descendants belongs to $\bm{Z}$.
Pearl (1993) focused on back-door paths, which can bias causal relationships, and established the back-door criterion. This criterion for covariate selection provides a sufficient condition for identifying total causal effects.
Definition 2.4 (Back-door criterion).
A set $\bm{Z}$ satisfying $\bm{Z}\cap\{X_{i},X_{j}\}=\emptyset$ is said to satisfy the back-door criterion relative to an ordered pair $(X_{j},X_{i})$ if
  1. (i)
    $\bm{Z}\cap\bm{DE}(X_{j})=\emptyset$, and
  2. (ii)
    $\bm{Z}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$.
Although the back-door criterion serves as a powerful framework for covariate selection, it suffers from a limitation in capturing partial causal effects originating from specific paths. To address this issue, we relax the first condition of the back-door criterion.
Definition 2.5 (Selective-door criterion).
A set $\bm{Z}$ satisfying $\bm{Z}\cap\{X_{i},X_{j}\}=\emptyset$ is said to satisfy the selective-door criterion relative to an ordered pair $(X_{j},X_{i})$ if
  1. (i)
    for any $X_{k}\in\bm{Z}\cap\bm{DE}(X_{j})$ such that there exists a directed path $X_{j}\rightarrow\rightarrow X_{k}$ that is opened by $\bm{Z}\setminus\{X_{k}\}$, every back-door path $X_{k}\dashleftarrow X_{i}$ is blocked by $(\bm{Z}\cup\{X_{j}\})\setminus\{X_{k}\}$, and
  2. (ii)
    $\bm{Z}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$.
Since the first condition (i) is weaker than that of the back-door criterion, the selective-door criterion allows some directed paths from $X_{j}$ to $X_{i}$ to be blocked. It should be noted, however, that blocking every path from $X_{j}$ to $X_{i}$ except for the directed paths of interest is not sufficient.
Figure 2: An example of unintended post-treatment bias induced by a non-collider
In Figure 2, the set $\{M_{2}\}$ only opens the directed path $X\rightarrow M_{1}\rightarrow Y$. However, this set does not satisfy the selective-door criterion relative to $(X,Y)$ due to the existence of the sequence of vertices $X\rightarrow M_{1}\rightarrow M_{2}\leftarrow M_{1}\rightarrow Y$. Since this structure is not a path and $M_{2}$ is not a collider, it can be easily overlooked when relying solely on the conventional concept of path blocking.
2.4 Controlled Total Effects
In the literature on linear SEMs, the decomposition of causal effects has been extensively discussed and forms the foundation of mediation analysis. In this framework, the direct effect of $X_{j}$ on $X_{i}$ is given by the path coefficient $a_{ij}$, which corresponds to the directed edge $X_{j}\to X_{i}$. The indirect effect of $X_{j}$ on $X_{i}$ is defined as the sum of products of path coefficients along each directed path $X_{j}\rightarrow\rightarrow X_{i}$, excluding the direct edge $X_{j}\to X_{i}$. The total effect of $X_{j}$ on $X_{i}$ is the sum of the direct and indirect effects. See Alwin and Hauser (1975) and Bollen (1987) for details on these causal effects.
To clarify the interpretation of partial regression coefficients, we define the following partial causal effects.
Definition 2.6 (Controlled total effect).
Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For a vertex set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$, consider the subgraph $G^{\prime}$ obtained by removing all edges whose terminal vertices belong to $\bm{Z}$. Then, the $\bm{Z}$-controlled total effect of $X_{j}$ on $X_{i}$, denoted by $\tau_{ij\mid do(\bm{Z})}$, is defined as the total effect of $X_{j}$ on $X_{i}$ restricted to the subgraph $G^{\prime}$.
The $\bm{Z}$-controlled total effect is the sum of all effects not mediated by variables in $\bm{Z}$. The validity of this measure is justified by the following alternative expression of the data-generating process, which is derived solely through successive substitution operations in the linear structural equations.
Proposition 2.7 (Ancestral expansion).
Assume that (ii) in Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). Then, for any subset $\bm{S}\subset\bm{V}\setminus\{X_{i}\}$, the following holds:
$X_{i}=\sum_{X_{j}\in\bm{S}}\tau_{ij\mid do(\bm{S}\setminus\{X_{j}\})}X_{j}+\sum_{X_{k}\in\bm{AN}(X_{i})\setminus\bm{S}}\tau_{ik\mid do(\bm{S})}u_{k}+u_{i}.$ (2.3)
If none of the omitted variables that constitute the error terms $u_{k}$ and $u_{i}$ are affected by the intervention on $X_{j}$, the following equality is justified:
$\tau_{ij\mid do(\bm{Z})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}=\bm{z}),\,do(X_{j}=x_{j})],$ (2.4)
where $\bm{Z}=\bm{S}\setminus\{X_{j}\}$, and “dodo” denotes the do-operator (Pearl, 2009). Note that the notation involving the do-operator denotes probabilistic concepts of the post-intervention distribution, i.e., the distribution obtained by replacing the right-hand side of the corresponding structural equation with a constant.
3 Causation and Regression
3.1 Problem Formulation
Suppose we are interested in the causal effect of $X_{j}$ on $X_{i}$. Let $\bm{S}\subset\bm{V}\setminus\{X_{i}\}$ denote the set of explanatory variables, and consider the following regression equation at the population level:
$X_{i}=\beta_{i0\mid\bm{S}}+\beta_{ij\mid\bm{Z}}X_{j}+\sum_{X_{k}\in\bm{Z}}\beta_{ik\mid\bm{S}\setminus\{X_{k}\}}X_{k}+\epsilon_{i\mid\bm{S}},$ (3.1)
where $\bm{Z}=\bm{S}\setminus\{X_{j}\}$ denotes a set of covariates. In this paper, the partial regression coefficient is defined using ordinary covariances, not conditional covariances. See the Appendix for details. The objective is to derive a sufficient condition on $\bm{Z}$ under which $\beta_{ij\mid\bm{Z}}=\tau_{ij\mid do(\bm{Z})}$ holds. If the coefficient $\beta_{ij\mid\bm{Z}}$ corresponds to the causal effect, then, under standard regularity conditions for random explanatory variables, the least squares estimator provides a consistent estimate of this effect.
3.2 Main Results
The selective-door criterion provides a sufficient condition of interest.
Theorem 3.1 (Selective-door adjustment).
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ satisfying the selective-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the $\bm{Z}$-controlled total effect $\tau_{ij\mid do(\bm{Z})}$.
In the case of linear regression, condition (i) of the selective-door criterion requires that if an explanatory variable $X_{k}\in\bm{Z}$ is a post-treatment variable of $X_{j}$, then one of the following conditions must hold: (1) all directed paths from $X_{j}$ to $X_{k}$ are blocked by other explanatory variables; or (2) all back-door paths from $X_{k}$ to $X_{i}$ are blocked by explanatory variables other than $X_{k}$. As will be shown later, the converse of Theorem 3.1 also holds almost everywhere in the parameter space. That is, in a generic sense, the selective-door criterion constitutes a necessary and sufficient condition for providing a linear causal interpretation of partial regression coefficients.
As corollaries, we obtain the following propositions.
Corollary 3.2 (D-separation adjustment).
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ that blocks every path between $X_{j}$ and $X_{i}$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) is equal to 0.
Corollary 3.3 (Single-door adjustment).
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{i}))$ that blocks every path between $X_{j}$ and $X_{i}$ except for the directed edge $X_{j}\rightarrow X_{i}$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the direct effect of $X_{j}$ on $X_{i}$.
Corollary 3.4 (Back-door adjustment).
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{j}))$ satisfying the back-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the total effect of $X_{j}$ on $X_{i}$.
These corollaries were originally established in Spirtes et al. (1998) and Pearl (1998). Note that the single-door criterion (adjustment) excludes descendants of the outcome variable. The reason for this exclusion is explained by the mechanism of post-treatment bias discussed in the following subsection.
In many empirical analyses, it is not uncommon for all explanatory variables to be regarded as causes of interest. Thus, we also provide sufficient conditions under which all regression coefficients can be interpreted as causal effects.
Proposition 3.5.
The following statements are equivalent:
  • (i)
    For any $X_{j}\in\bm{S}$, the set $\bm{S}\setminus\{X_{j}\}$ satisfies the selective-door criterion relative to $(X_{j},X_{i})$.
  • (ii)
    For any $X_{j}\in\bm{S}$, the set $\bm{S}\setminus\{X_{j}\}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$.
  • (iii)
    For any $X_{j}\in\bm{S}$, the set $\bm{S}\setminus\{X_{j}\}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$ that contains no v-structure.
Condition $(iii)$ is particularly easy to check because $\bm{S}\setminus\{X_{j}\}$ blocks a back-door path from $X_{j}$ to $X_{i}$ that contains no v-structure if and only if $X_{i}\notin\bm{S}$ and at least one vertex $X_{k}\in\bm{S}\setminus\{X_{j}\}$ lies on that back-door path.
3.3 Post-Treatment Bias
Biases that arise from conditioning on post-treatment variables (i.e., descendants in the causal diagram) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the selective-door criterion is required.
In addition to the set $\bm{S}=\{X_{j}\}\cup\bm{Z}$ in (3.1), the following disjoint subsets are introduced:
$\bm{S}_{1}$ $=\{\,X_{k}\in\bm{S}\mid\text{$\bm{S}\setminus\{X_{k}\}$ opens a back-door path $X_{k}\dashleftarrow X_{i}$}.\,\}$
$\bm{S}_{2}$ $=\{\,X_{k}\in\bm{S}\mid\text{$\bm{S}\setminus\{X_{k}\}$ blocks every back-door path $X_{k}\dashleftarrow X_{i}$}.\,\}$
The bias $\gamma_{ik\mid\bm{S}\setminus\{X_{k}\}}$ corresponding to each coefficient $\beta_{ik\mid\bm{S}\setminus\{X_{k}\}}$ in (3.1) is defined as
$\gamma_{ik\mid\bm{S}\setminus\{X_{k}\}}\coloneq\beta_{ik\mid\bm{S}\setminus\{X_{k}\}}-\tau_{ik\mid do(\bm{S}\setminus\{X_{k}\})},\quad X_{k}\in\bm{S}.$
Although bias is often defined as the difference from the total effect, we adopt the above definition for clarity in the quantitative characterization of post-treatment bias.
The following theorem characterizes post-treatment bias while controlling for confounding bias.
Theorem 3.6.
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ that blocks every back-door path $X_{j}\dashleftarrow X_{i}$, the bias of the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) is given by
$\gamma_{ij\mid\bm{Z}}=-\sum_{X_{p}\in\bm{S}_{1}}\gamma_{ip\mid\bm{S}\setminus\{X_{p}\}}\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}.$
As indicated by the back-door criterion, confounding bias can be eliminated by blocking every back-door path. The first condition (i) of the selective-door criterion implies that $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ and can be understood as a condition to prevent post-treatment bias. In particular, conditioning on a descendant $X_{p}$ of the outcome variable $X_{i}$ is problematic, since whenever there exists an opened directed path $X_{j}\rightarrow\rightarrow X_{i}$, a sequence of vertices $X_{j}\rightarrow\rightarrow X_{i}\rightarrow\rightarrow X_{p}\leftarrow\leftarrow X_{i}$ exists, leading to post-treatment bias. On the other hand, if each path from $X_{j}$ to $X_{i}$ is blocked (i.e., $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ ), the partial regression coefficient vanishes even when a decendant of the outcome variable $X_{i}$ is included among the conditioning variables.
3.4 Necessity for Almost All Parameter Points
Here, we discuss the precise scope of applicability of the selective-door criterion. For the coefficient matrix $\bm{A}$ and covariance matrix $\bm{\Sigma}$, let $(\bm{A},\bm{\Sigma})$ denote a parameter of the model (2.1). Since covariance information is sufficient to characterize the partial regression coefficients of interest, we omit the intercept term $\bm{c}$ from the model parameter. We consider the following parameter space:
$\bm{\Theta}\coloneq\{(\bm{A},\bm{\Sigma})\mid\text{ $\bm{A}$ and $\bm{\Sigma}$ satisfy Assumption (2.1)}\}.$ (3.2)
Almost everywhere in this parameter space, satisfying the selective-door criterion is not only sufficient but also necessary for partial regression coefficients to admit a linear causal interpretation.
Theorem 3.7.
Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). Fix a set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ and consider the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1). Then, except on a set of Lebesgue measure zero in the parameter space $\bm{\Theta}$, the following statements are equivalent:
  1. (i)
    The set $\bm{Z}$ satisfies the selective-door criterion relative to $(X_{j},X_{i})$.
  2. (ii)
    The coefficient $\beta_{ij\mid\bm{Z}}$ coincides with the $\bm{Z}$-controlled total effect $\tau_{ij\mid do(\bm{Z})}$.
3.5 Application to Nonlinear Data-Generating Processes
Since simple linear models have limitations in their approximation performance, we consider the following structural equation model:
$X_{i}=\sum_{X_{j}\in\bm{PA}(X_{i})}a_{ij}X_{j}+\sum_{k}a_{h_{ik}}h_{ik}(\bm{PA}(X_{i}))+u_{i},$ (3.3)
where each $h_{ik}$ is a known nonlinear function of $\bm{PA}(X_{i})$, and each term $h_{ik}(\bm{PA}(X_{i}))$ can capture interactions and effect modification (see VanderWeele (2009)). This model accommodates spline-based models (Cox, 1972; de Boor, 1972; Wahba, 1978; Silverman, 1985; Eilers and Marx, 1996). Even when nonlinear terms are present, the discussion so far remains applicable by explicitly representing them as vertices in the graph and replacing the directed edges corresponding to nonlinear causal relationships with bidirected edges. In the following, for the sake of clarity, we illustrate this point through a concrete example.
To accommodate nonlinear structures, we impose an additional standard assumption on the causal path diagram.
Assumption 3.8.
Any random vector consisting only of error terms corresponding to vertices that are not adjacent to $X_{i}$ by a bidirected edge is independent of $u_{i}$.
$\begin{cases}Z=u_{Z}\\ X=a_{XZ}Z+u_{X}\\ M=a_{MX}X+a_{MZ}Z+u_{M}\\ H=h(X,M)=u_{H}\\ Y=a_{YX}X+a_{YM}M+a_{YH}H+a_{YZ}Z+u_{Y},\end{cases}$ (3.4)
Figure 3: The causal path diagram associated with (3.4)
Assumption 3.8 guarantees the absence of correlation between $u_{H}$ and $u_{Y}$ when $u_{Z}$, $u_{X}$, and $u_{M}$ are independent of $u_{Y}$. Depending on the nonlinear function $h$, some of the bidirected edges in Figure 3 may be omitted, since uncorrelatedness may hold. This causal path diagram shows that a bidirected edge does not necessarily indicate the presence of a confounding factor; it may also represent a causal relationship that cannot be expressed by path coefficients. For instance, the paths $X\leftrightarrow H\to Y$ and $X\to M\leftrightarrow H\to Y$ mediate the nonlinear causal effect of $X$ on $Y$. Since model (3.3) can be described as a special case of model (2.1), as in model (3.4), all the results obtained thus far are applicable to the identification of linear causal effects.
In this case, the following decomposition of the “true” total effect holds:
$\frac{d}{dx}\,\mathrm{E}[Y\mid do(X=x)]$ $=\tau_{YX}+a_{YH}\frac{d}{dx}\,\mathrm{E}[h(X,M)\mid do(X=x)],$
where $\tau_{YX}=a_{YX}+a_{YM}a_{MX}$ is the total effect in the sense of linear SEM. More generally, this decomposition is expressed as
$\tau_{YX}\coloneq\frac{\partial}{\partial x}\,\mathrm{E}[Y\mid do(\bm{H}=\bm{h}),\,do(X=x)],$
$\delta_{YX}(x)\coloneq\frac{d}{dx}\,\mathrm{E}[Y\mid do(X=x)]-\tau_{YX},$
$\frac{d}{dx}\,\mathrm{E}[Y\mid do(X=x)]=\tau_{YX}+\delta_{YX}(x),$
where $\bm{H}$ denotes all nonlinear terms. The second term $\delta_{YX}(x)$ represents the causal effects originated from nonlinear structures and includes the mediated interactive effect (VanderWeele, 2013). If the nonlinear terms are not differentiable, one can instead consider the difference between two intervention levels, $do(X=x)$ and $do(X=x^{\prime})$. In general, the identifiability of the nonlinear causal effect $\delta_{YX}(x)$ depends on the form of nonlinear functions and causal relationships among the ancestors of $Y$.
Here, we restrict our discussion to the case of (3.4). Note that the derivative constituting the nonlinear effect can be expressed as follows.
$\frac{d}{dx}\,\mathrm{E}[h(X,M)\mid do(X=x)]=\frac{d}{dx}\,\mathrm{E}\left[\tilde{h}\left(x,a_{MX},a_{MZ}Z+u_{M}\right)\right],$ (3.5)
where $\tilde{h}$ is a function. Fortunately, this nonlinear causal effect of $X$ on $H$ can be identified because $a_{MX}=\beta_{MX\mid Z}$ and $a_{MZ}Z+u_{M}=M-\beta_{MX\mid Z}X$ hold in the following regression equation:
$M=\beta_{M0\mid XZ}+\beta_{MX\mid Z}X+\beta_{MZ\mid X}Z+\epsilon_{M\mid XZ}.$
Moreover, $\tau_{YX}=\beta_{YX\mid HZ}$ and $a_{YH}=\beta_{YH\mid XZ}$ hold in the following regression equation:
$Y=\beta_{Y0\mid XHZ}+\beta_{YX\mid HZ}X+\beta_{YH\mid XZ}H+\beta_{YZ\mid XH}Z+\epsilon_{Y\mid XHZ}.$
Therefore, the effects $\tau_{YX}$, $\delta_{YX}(x)$ and $\tau_{YX}+\delta_{YX}(x)$ are identifiable if we can also observe $M$ and $Z$.
4 Discussion
We have derived generalized theorems that encompass existing results characterizing partial regression coefficients based on conventional graphical criteria. In causal diagrams represented by ADMGs, bidirected edges indicate the presence of unobserved common causes. In contrast, in the causal path diagrams defined in this paper, bidirected edges represent only correlations among the error terms, which constitutes a major difference. As in the case of nonlinear terms, for graphs with cycles or for partial ancestral graphs (Zhang, 2008), the results of this paper remain applicable as long as the graphs can be transformed into an ADMG while disregarding the usual interpretation of bidirected edges. This is largely due to the fact that the all proofs of the theorems in this paper rely only on algebraic arguments.
In the standard linear SEM, where bidirected edges represent the presence of unobserved confounders, the linear causal interpretation of partial regression coefficients under Theorem 3.1 can be summarized as follows:
$\beta_{ij\mid\bm{Z}}$ $=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j}),\,do(\bm{Z}^{j}=\bm{z}^{j})]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j}),\,\bm{Z}^{j}=\bm{z}^{j}]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j})],$
where $\bm{Z}_{j}\coloneq\bm{Z}\cap\bm{DE}(X_{j})$, and $\bm{Z}^{j}\coloneq\bm{Z}\setminus\bm{Z}_{j}$. By definition, it is clear that $\tau_{ij\mid do(\bm{Z})}\equiv\tau_{ij\mid do(\bm{Z}_{j})}$. In particular, when $\bm{Z}_{j}\neq\emptyset$, special attention must be paid to the following relationship.
$\beta_{ij\mid\bm{Z}}$ $\not\equiv\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid\bm{Z}=\bm{z};\,do(X_{j}=x_{j})]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(X_{j}=x_{j}),\,\bm{Z}=\bm{z}].$
Here, $\mid\bm{Z}=\bm{z};\,do(X_{j}=x_{j})$ denotes intervening on $X_{j}$ after observing $\bm{Z}=\bm{z}$, and is conceptually similar to the counterfactual query $Q_{c}$ (Pearl, 2015). If the back-door criterion is satisfied, the symbol $\not\equiv$ can be replaced with ==. These interpretations rely on the absence of interaction or effect modification, and therefore does not necessarily hold when nonlinear terms are present. See the Appendix for details.
As a direction for future research, a more detailed investigation of the estimation of decomposed nonlinear causal effects is required. In particular, the discussion of identifiability of nonlinear causal effects in this paper is limited to whether the causal parameter (or its function) can be expressed as a function of the observed distribution. For example, the interchangeability of differentiation and expectation in (3.5) is expected to affect estimation accuracy. Therefore, further analysis under appropriate assumptions on the nonlinear terms remains an important topic for future work.
Acknowledgments
The author is grateful to Professor Kou Fujimori for helpful comments and encouragement during the preparation of this manuscript.
Statements and Declarations
The author has no relevant financial or non-financial interests to disclose.